OpenStudy (anonymous):

Find the slope and the y-intercept of the line. 4y= 3x-20

hero (hero):

divide both sides by 4 to get... y = .75x - 5 Slope = .75 or 3/4 y-int = -5

OpenStudy (anonymous):

Y intercept by definition is a point on the y-axis, i.e. where x coordinate is 0. put x = 0 to get the y intercept: y = -20/4 = -5. Subsequently, slope: which is by definition the change in y coordinate for a unit change in x coordinate. So what you could try is move a point from x = 0 to x = 1 (which is a unit change) and note the change in y. Now, at x = 1 y = -17/4, i.e. it changed from -5 to -17/4 or net change in y = 3/4 (which would be the slope)